OpenStudy (anonymous):

Write an equation for the line that contains the point (-6, -8) and is parallel to the line 3x – 2y = -12. A. -3x + 2y = -10 B. -3x + 2y = 2 C. -3x + 2y = 22 D. 2x + 3y = -36

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

i think its b for some reason

OpenStudy (anonymous):

Solve 3x – 2y = -12 for y.\[y=\frac{3 x}{2}+6 \]Solve the following for y:\[\frac{y--8}{x--6}=\frac{3}{2} \]

OpenStudy (anonymous):

im sorry but there are ????all over what youwrote

OpenStudy (anonymous):

b. is the answer.

OpenStudy (anonymous):

okthank you so much

OpenStudy (anonymous):

can i ask 1 more?

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

Refer to the attachment

OpenStudy (anonymous):

ok well this is a different question

OpenStudy (dtan5457):

The answer is indeed B. How you do it if you want the explanation is this. Given the equation 3x – 2y = -12. You can simplfy that equation to -2y=-12-3x divide that by -2. y=6+3/2x 3/2 is your slope point slope formula y-y1=M(x-x1) your point is -6,-8 y+8=3/2(x+6) simplify that to y+8=3/2x+9 that's the same as y=3/2x+1 all of your answer choices have 2y. so multiply this by and you get 2y=3x+2 you can shape that to -3x + 2y = 2 plug that in to y

OpenStudy (anonymous):

thank you @dtan5457

OpenStudy (dtan5457):

I answered your latest question, go check it out

OpenStudy (anonymous):

i did thanks again